Nitrogen and hydrogen can react to form ammonia.
N2(g)+3H2(g)→2NH3(g)

1. Choose ratios showing the relationships between moles of each of the reactants and products in the reaction. Multiple choice, which one?

A. 2 mol N2 : 2 mol H2 : 3 mol NH3

B.1 mol N2 : 1/3 mol H2 : 1/2 mol NH3

C. 1 mol N2 : 1 mol H2 : 1 mol NH3

D. 1 mol N2 : 3 mol H2 : 2 mol NH3

2. How many molecules of H2 are required to completely react with two molecules of N2?

Express your answer as an integer.

3. How many moles of H2 are required to completely react with 2 mol of N2?

Express your answer using one significant figures.
Answer #1

1) option D is correct.

number of moles of N2 are 1 number of moles of h2 are 3 and the number of moles of NH3 are 2.

2)

6.023*1023 molecules of N2 = 1 mole of n2 so

2 molecules of N2 = 2 molecules of N2*(1 mole of n2 /6.023*1023 molecules of N2)

= 0.332*10-23 mole of N2

1 mole of N2 requires 3 mole of H2 so

0.332*10-23 mole of N2 will require 0.996*10-23 mole of H2

1 mole of H2 = 6.023*10^23 molecules so

0.996*10-23 mole of H2 = 0.996*10-23 mole of H2*(6.023*10^23 molecules/1 mole of H2)

= 6 molecules of h2

Therefore,the number of molecules of h2 required will be 6.

3)

from the balanced equation we can say that

1 mole of N2 requires 3 mole of H2 so

2 mole of N2 will require

= 2 mole of N2 *(3 mole of H2/1 mole of N2)

= 6 mole of H2

Therefore, the number of moles of H2 required will be 6